# Learning Individual Reproductive Behavior from Aggregate Fertility Rates via Neural Posterior Estimation

### v1

@@ Abstract
While age-specific fertility rates (ASFRs) provide the most extensive record of reproductive change, their aggregate nature masks the underlying behavioral mechanisms that ultimately drive fertility trends. To recover these mechanisms, we develop a likelihood-free Bayesian framework that couples an individual-level model of the reproductive process with Sequential Neural Posterior Estimation (SNPE). This allows us to infer eight behavioral and biological parameters from just two aggregate series: ASFRs and the age-profile of planned versus unplanned births. Applied to U.S. National Survey of Family Growth cohorts and to Demographic and Health Survey cohorts from Colombia, the Dominican Republic, and Peru, the method reproduces observed fertility schedules and, critically, predicts out-of-sample micro-level distributions of age at first sex, inter-birth intervals, and family-size ideals, none of which inform the estimation step. Because the fitted model yields complete synthetic life histories, it enables behaviorally explicit population forecasts and supports the construction of demographic digital twins.
@@ 1 Introduction
Globally declining fertility rates are driving a profound demographic transformation, presenting first-order societal challenges worldwide and underscoring the critical need for robust models to understand and anticipate future fertility trends. Age-specific fertility rates (ASFRs) —births per woman at each age— remain the workhorse input for these tasks because the underlying birth counts are collected continuously in almost every country’s vital-statistics system. ASFRs series therefore provide unparalleled temporal and geographic coverage.
Yet ASFRs are only marginal aggregates. They reveal when births occur but not why. A long tradition of parametric models has been developed to describe ASFRs curves, smooth observed schedules, extrapolate fitted parameters, or convert projected total fertility rates into age patterns for cohort-component population forecasts (Coale and Trussell, 1974; Brass, 1974; Schmertmann, 2003; Chandola et al., 1999). While existing approaches can excel at capturing aggregate ASFRs shapes, their parameters are typically macro-level abstractions. As such, they offer limited or elusive insight into the individual behavioural mechanisms driving fertility decisions (Hoem and Hoem, 1989). Consequently, forecasts based on these models often cannot explicitly account for how evolving individual choices regarding the timing of parenthood, contraceptive practice, and birth-spacing and stopping decisions shape future fertility trends.
@@
This paper proposes a step toward this goal. We introduce a demographically interpretable individual-level stochastic model that captures key features of modern reproductive behavior, including stopping, spacing, and contraceptive failure. Estimating such a richly parameterised model from coarse data is a non-trivial inverse problem: many combinations of micro parameters can generate similar ASFRs curves. We show that adding the age profile of planned versus unplanned births to the set of ASFRs, provides enough information to identify the key parameters of this model. We fit the model to U.S. National Survey of Family Growth cohorts and to Demographic and Health Survey cohorts from Colombia, the Dominican Republic, and Peru, representing four settings that span markedly different fertility regimes.
Using Sequential Neural Posterior Estimation (SNPE), a likelihood-free method that learns to approximate posterior parameter distributions from simulator outputs, we recover full posterior distributions of behaviourally meaningful parameters. Posterior predictive checks further confirm that the inferred parameters enable the simulator to accurately reproduce the observed aggregate fertility and unplanned fertility rates. Most significantly, we conduct out-of-sample validation, demonstrating that the parameters inferred solely from these aggregate population rates can successfully predict distributions of individual-level characteristics, such as the empirical distributions of age at first sexual intercourse, desired family sizes reported in surveys, and realized birth intervals. This capacity to predict micro-level outcomes not directly used in the estimation process provides strong evidence for the model’s structural validity and the inference procedure’s power.
By establishing a demographically interpretable and statistically robust link between individual reproductive behaviors and population-level fertility outcomes, our framework offers a promising path towards more nuanced understanding and potentially more accurate and behaviorally grounded fertility forecasts.
@@ 1.1 Previous Work
Our own previous work showed that such a bridge is feasible in natural-fertility contexts —historical or religious communities without deliberate birth control— where a handful of biological parameters govern the reproductive process and can be inferred from ASFRs alone (Ciganda and Todd, 2024). The present paper tackles the far more ambitious, and policy-relevant task, of modelling contemporary fertility, where stopping, spacing, and imperfect contraception introduce complex decision making and heterogeneity in reproductive dynamics.
@@ 1.2 Organization of the Article
The remainder of this paper is structured as follows: Section 2 describes our individual-level computational model of the reproductive process. Section 3 details the data sources and the construction of ASFRs and Age-specific unplanned fertility rates (ASUFRs). Section 4 outlines the SNPE methodology. The Results section presents our findings on parameter identifiability via cross-validation, model fit through posterior predictive checks, and the out-of-sample validation against micro-level data. Finally, we discuss the implications of our findings, limitations, and avenues for future research in the Discussion.
@@ 2.1 Implementation
The dynamics described above are implemented by assigning the following characteristics to each woman $i$ and defining the processes that govern her reproductive journey:
@@
    Age at Sexual Initiation ($X_{i}$): $X_{i}$ (in months) is drawn from a lognormal distribution. This distribution is parameterized by an estimated mean age at sexual initiation $\mu_{s}$ (in years) and an estimated standard deviation $\sigma_{s}$ (in years). These are converted to the lognormal distribution’s underlying parameters (mean of logarithms $\mu_{\ln,s}$ and standard deviation of logarithms $\sigma_{\ln,s}$) using the standard formulas: $\mu_{\ln,s}=\ln(\mu_{s}^{2}/\sqrt{\mu_{s}^{2}+\sigma_{s}^{2}})$ and $\sigma_{\ln,s}=\sqrt{\ln(1+\sigma_{s}^{2}/\mu_{s}^{2})}$. Ages are then drawn and converted to months.
@@
    Age at Intentional Reproduction ($R_{i}$): $R_{i}$ (in months) is also drawn from a lognormal distribution. The mean of this distribution is set to $\mu_{r}=\mu_{s}+\delta_{r}$, where $\delta_{r}$ (in years) is an estimated model parameter. The standard deviation for the age at intentional reproduction, $\sigma_{r}$, is fixed at 2.6 years. The parameters $\mu_{r}$ and $\sigma_{r}$ (in years) are used to derive the underlying lognormal parameters $\mu_{\ln,r}$ and $\sigma_{\ln,r}$ using the same conversion formulas.
@@
    Desired Family Size ($D_{i}$): $D_{i}$ is drawn from a Weibull distribution, parameterized by an estimated mean desired family size $\mu_{d}$ and a fixed standard deviation $\sigma_{d}=1.25$ children. These are used to approximate the Weibull shape parameter $\alpha_{d}=(\sigma_{d}/\mu_{d})^{-1.086}$ and scale parameter $\lambda_{d}=\mu_{d}/\Gamma(1+1/\alpha_{d})$. Samples are rounded to the nearest integer.
@@
    Desired Birth Spacing ($B_{i}$): $B_{i}$ (in months) is drawn from a lognormal distribution, parameterized by an estimated mean desired spacing $\mu_{b}$ (in months) and standard deviation $\sigma_{b}=27$ months. The underlying lognormal parameters $\mu_{\ln,b}$ and $\sigma_{\ln,b}$ are derived similarly.
We fix the dispersion parameters ($\sigma_{r},\sigma_{d},\sigma_{b}$) at plausible, empirically-informed values because, while contributing to overall individual heterogeneity, these second-order moments were found to exert a less direct or distinguishable influence on the aggregate ASFRs and ASUFR patterns compared to parameters governing central tendencies or the core dynamics of conception and contraception. Furthermore, robustly identifying these specific dispersion parameters solely from the available aggregate fertility schedules proved challenging in preliminary analyses. Fixing them allows the inferential procedure to concentrate more effectively on the primary behavioral and physiological parameters that are both more influential at the aggregate level and better identified by the data, without excessively constraining the model’s ability to capture essential individual-level variability.
@@
Upon simulating complete reproductive trajectories, we derive aggregate age-specific fertility rates (ASFRs). However, relying solely on ASFRs for parameter inference may be insufficient given the depth of behavioral detail incorporated in our model. Without distinguishing births that align with stated intentions from those that are mistimed or unwanted, different parameter configurations (e.g., high desired family size with effective contraception versus lower desired family size with poor contraception) could yield deceptively similar ASFRs, creating critical identifiability challenges. Therefore, alongside ASFRs, a key output of our model is a vector of age-specific unplanned fertility rates (ASUFRs). These ASUFRs, detailed further in Section 3, exclusively take into account births classified as mistimed or unwanted, providing essential information to better constrain the model’s parameters and narrow the range of configurations consistent with the observed data.
Table 1 summarizes the key estimated model parameters.
@@
$\mu_{s}$ Mean age at sexual initiation (in years), modeled with a lognormal distribution.
$\sigma_{s}$ Standard deviation for the distribution of ages at sexual initiation.
@@
$\mu_{d}$ Mean desired family size, modeled using a Weibull distribution.
$\kappa$ Probability of contraceptive failure.
$\mu_{b}$ Mean desired birth spacing (in months), modeled with a lognormal distribution.
@@
Table 1: Key model parameters and their descriptions.
@@ 3 Data
Our analysis draws on two large, nationally representative sources. For countries outside the United States, we use the Demographic and Health Surveys (DHS), a leading initiative of the United States Agency for International Development (USAID) that has conducted over 300 surveys in more than 90 low-and middle-income countries. Each DHS round collects complete retrospective birth histories, allowing the reconstruction of every woman’s fertility trajectory. Data for the United States are drawn from several early cycles of the National Survey of Family Growth (NSFG), specifically the 1972, 1976, 1982, and 1988 rounds. These surveys, conducted by the National Center for Health Statistics (NCHS), are designed to provide detailed national data on family life, marriage, divorce, contraception, and importantly for this study, complete pregnancy and birth histories. We utilize these data as compiled and harmonized by the Integrated Fertility Survey Series (IFSS), applying IFSS sampling weights to generate population-level estimates. For both DHS and NSFG data, to achieve adequate sample sizes while ensuring cohorts of women were exposed to broadly similar social conditions, we pool available survey rounds and organize respondents into 10-year birth cohorts.
@@
Because DHS sample sizes and the number of survey rounds differ widely, we implemented a further filter based on cohort size. We selected DHS countries where the chosen cohort included a minimum of 4,000 women, a threshold established to guarantee reasonably precise estimates of both age-specific total and unplanned fertility rates across the entire reproductive span. This resulted in the inclusion of three DHS countries: Colombia, Peru, and the Dominican Republic. Together with the United States, these four nations constitute our study sample, offering a valuable cross-section of different fertility patterns and socio-economic contexts that allows for a robust examination of our model’s performance across diverse demographic and social settings.
We compute the conventional age-specific fertility rate (ASFRs) by dividing the number of births at each age by the number of women in the cohort at that age. To construct the age-specific unplanned fertility rate (ASUFR) we combine each birth’s intention status with the respondent’s ideal family size. In the DHS, questions about birth intention (i.e., whether a birth was “wanted then,” “wanted later,” or “not wanted”) are asked only for events in the 3–5 years preceding the interview. We therefore estimate, for every maternal age, the share of below-ideal births classed as unplanned in that recent window and extrapolate it to earlier events, while always counting births that push parity above the ideal as unwanted. In the NSFG, by contrast, intention status is recorded for all births, so we classify each event directly from the survey response and apply the same parity rule for above-ideal births. By coupling direct reports where available with an indirect parity check where necessary, our strategy mitigates recall bias and ex-post rationalization, which are known to severely affect unplanned fertility estimates obtained from the intention questions alone (Casterline and El-Zeini, 2007).
@@ 4 Simulation-Based Inference
Inferring parameters for complex process-based models, such as our individual-level simulation of reproductive trajectories, is often challenging due to intractable likelihood functions. Neural Posterior Estimation (NPE) offers a powerful simulation-based inference (SBI) approach that circumvents these difficulties by directly learning an approximation to the posterior distribution $p(\theta\mid x)$ from simulated data pairs $(\theta,x)$. In practice, NPE proceeds by drawing parameters $\theta$ from a prior distribution ${p}(\theta)$, using the simulator to generate the corresponding data $x\sim p(x\mid\theta)$, and then training a flexible neural density estimator, $q_{\phi}(\theta\mid x)$ (with network parameters $\phi$), on these simulated pairs to approximate the true posterior.
While basic NPE is effective, our study employs a powerful sequential variant known as Automatic Posterior Transformation (APT), sometimes referred to as SNPE-C (Greenberg et al., 2019). Like other sequential neural posterior estimation (SNPE) methods, APT enhances inferential efficiency by drawing parameter samples from an iteratively updated proposal distribution, $\tilde{p}(\theta)$, rather than solely from the fixed prior. The key challenge in this approach is to correct for the use of a proposal distribution, as naively training on the resulting samples would lead to an incorrect posterior estimate. APT solves this by analytically transforming the posterior estimate itself. For each simulated data point drawn from a proposal $\tilde{p}_{r}(\theta)$, it re-weights the target density using the known ratio $\tilde{p}_{r}(\theta)/p(\theta)$. The network is then trained by minimizing the loss with respect to this transformed "proposal posterior" (see line 10 of Algorithm 1). A key advantage of this formulation is that because the loss is valid for any proposal, the network can be trained on the cumulative set of all simulations generated across all previous rounds. This procedure allows the network to learn the true posterior without relying on the potentially unstable importance weights used by other methods.

### v2

@@ Abstract
Age-specific fertility rates (ASFRs) provide the most extensive record of reproductive change, but their aggregate nature obscures the individual-level behavioral mechanisms that drive fertility trends. To bridge this micro-macro divide, we introduce a likelihood-free Bayesian framework that couples a demographically interpretable, individual-level simulation model of the reproductive process with Sequential Neural Posterior Estimation (SNPE). We show that this framework successfully recovers core behavioral parameters governing contemporary fertility, including preferences for family size, reproductive timing, and contraceptive failure, using only ASFRs. The framework’s effectiveness is validated on cohorts from four countries with diverse fertility regimes. Most compellingly, the model, estimated solely on aggregate data, successfully predicts out-of-sample distributions of individual-level outcomes, including age at first sex, desired family size, and birth intervals. Because our framework yields complete synthetic life histories, it significantly reduces the data requirements for building microsimulation models and enables behaviorally explicit demographic forecasts.
@@ 1 Introduction
Globally declining fertility rates are driving a profound demographic transformation, presenting first-order societal challenges and underscoring the need for robust models to understand and anticipate future fertility trends. Age-specific fertility rates (ASFRs), defined as the number of births to women of a given age per person-year of exposure, remain the workhorse input for these tasks because the underlying birth counts are collected continuously in every country’s vital-statistics system. ASFRs series therefore provide unparalleled temporal and geographic coverage.
Yet ASFRs are only marginal aggregates. They reveal when births occur but not why. A long tradition of parametric models has been developed to describe ASFRs curves, smooth observed schedules, extrapolate fitted parameters, or convert projected total fertility rates into age patterns for cohort-component population forecasts (Coale and Trussell, 1974; Brass, 1974; Schmertmann, 2003; Chandola et al., 1999). While existing approaches can excel at capturing aggregate ASFRs shapes, their parameters are typically macro-level abstractions. As such, they offer limited insight into the individual behavioural mechanisms driving fertility decisions (Hoem and Hoem, 1989). Consequently, forecasts based on these models often cannot explicitly account for how evolving individual choices regarding the timing of parenthood, contraceptive practice, and birth-spacing and stopping decisions shape future fertility trends.
@@
This paper proposes a step toward this goal. We introduce a demographically interpretable individual-level model that captures key features of modern reproductive behavior, including stopping, spacing, and contraceptive failure. Estimating such a richly parameterised model from coarse data is a significant challenge, as many combinations of micro-parameters can generate similar ASFRs curves. We overcome this challenge by pairing our individual-level, mechanistic model with Sequential Neural Posterior Estimation (SNPE), a powerful simulation-based inference method. This approach allows us to demonstrate that core behavioral parameters governing the overall level and timing of fertility can be recovered from ASFRs alone. We then fit the model to U.S. National Survey of Family Growth cohorts and to Demographic and Health Survey cohorts from Colombia, the Dominican Republic, and Peru, representing four settings that span markedly different fertility regimes.
The framework’s effectiveness is confirmed through rigorous validation. First, posterior predictive checks show that the observed aggregate fertility rates are highly plausible under the model, falling well within the 95% posterior predictive intervals. Most significantly, we conduct out-of-sample validation, demonstrating that the model can successfully predict distributions of individual-level characteristics, such as the empirical distributions of age at first sexual intercourse, desired family size, and birth intervals, none of which inform the estimation step.
By establishing a demographically interpretable and statistically robust link between individual reproductive behaviors and population-level fertility outcomes, our framework offers a promising path towards a more nuanced understanding of fertility dynamics. This approach not only enables more accurate, behaviorally grounded population forecasts but also lowers the barrier to building detailed microsimulation models by significantly reducing their data requirements.
@@ 1.1 Previous Work
Our own previous work showed that such a bridge is feasible in natural-fertility contexts (historical or religious communities without deliberate birth control) where a handful of biological parameters govern the reproductive process and can be inferred from ASFRs alone (Ciganda and Todd, 2024). The present paper tackles the far more ambitious and policy-relevant task of modeling contemporary fertility, where stopping, spacing, and imperfect contraception introduce complex decision-making and heterogeneity in reproductive dynamics.
@@ 1.2 Organization of the Article
The remainder of this paper is structured as follows: Section 2 describes our individual-level computational model of the reproductive process. Section 3 details the data sources and the construction of the main demographic quantities used in our analysis. Section 4 outlines the SNPE methodology. The Results section presents our findings on parameter identifiability via cross-validation, model fit through posterior predictive checks, and the out-of-sample validation against micro-level data. Finally, we discuss the implications of our findings, limitations, and avenues for future research in the Discussion.
@@ 2.1 Implementation
The dynamics described above are implemented by assigning the following characteristics to each woman $i$ and defining the processes that govern her reproductive trajectory:
@@
    Age at Sexual Initiation ($X_{i}$): Each woman’s age at sexual initiation, $X_{i}$, is drawn from a lognormal distribution. This distribution is parameterized by a mean age $\mu_{s}$ and a standard deviation $\sigma_{s}$ (both in years), which are converted to the distribution’s underlying parameters (the mean of logarithms, $\mu_{\ln,s}$, and the standard deviation of logarithms, $\sigma_{\ln,s}$) using the standard formulas: $\mu_{\ln,s}=\ln(\mu_{s}^{2}/\sqrt{\mu_{s}^{2}+\sigma_{s}^{2}})$ and $\sigma_{\ln,s}=\sqrt{\ln(1+\sigma_{s}^{2}/\mu_{s}^{2})}$. The value for each woman is drawn in years and then converted to months.
@@
    Age at Intentional Reproduction ($R_{i}$): $R_{i}$ (in months) is also drawn from a lognormal distribution. The mean of this distribution is set to $\mu_{r}=\mu_{s}+\delta_{r}$, where $\delta_{r}$ (in years) is an estimated model parameter. The parameters $\mu_{r}$ and $\sigma_{r}$ (in years) are used to derive the underlying lognormal parameters $\mu_{\ln,r}$ and $\sigma_{\ln,r}$ using the same conversion formulas.
@@
    Desired Family Size ($D_{i}$): $D_{i}$ is drawn from a Weibull distribution, parameterized by its mean $\mu_{d}$ and standard deviation $\sigma_{d}$, which are used to approximate the Weibull shape parameter $\alpha_{d}=(\sigma_{d}/\mu_{d})^{-1.086}$ and scale parameter $\lambda_{d}=\mu_{d}/\Gamma(1+1/\alpha_{d})$. Samples are rounded to the nearest integer.
@@
    Desired Birth Spacing ($B_{i}$): $B_{i}$ (in months) is drawn from a lognormal distribution with a mean desired spacing $\mu_{b}$ and a standard deviation $\sigma_{b}$ (both in months). These parameters are used to derive the underlying mean and standard deviation for the lognormal distribution.
@@
Table 1 summarizes the estimated model parameters.
@@
$\mu_{s}$ Mean age at sexual initiation (in years), for a lognormal distribution.
$\sigma_{s}$ Standard deviation of the age at sexual initiation.
@@
$\sigma_{r}$ Standard deviation of the age at start of intentional reproduction.
$\mu_{d}$ Mean desired family size, for a Weibull distribution.
$\sigma_{d}$ Standard deviation of desired family size.
$\mu_{b}$ Mean desired birth spacing (in months), for a lognormal distribution.
$\sigma_{b}$ Standard deviation of the desired birth spacing.
$\kappa$ Probability of contraceptive failure per month.
@@
Table 1: Model parameters and their descriptions.
@@ 3 Data
Our analysis draws on two large, nationally representative sources. For countries outside the United States, we use the Demographic and Health Surveys (DHS), a leading initiative of the United States Agency for International Development (USAID) that has conducted over 300 surveys in more than 90 low-and middle-income countries. Each DHS round collects complete retrospective birth histories, allowing the reconstruction of every woman’s fertility trajectory. Data for the United States are drawn from several early cycles of the National Survey of Family Growth (NSFG). These surveys, conducted by the National Center for Health Statistics (NCHS), are designed to provide detailed national data on family life, marriage, divorce, contraception, and importantly for this study, complete pregnancy and birth histories. We utilize these data as compiled and harmonized by the Integrated Fertility Survey Series (IFSS), applying IFSS sampling weights to generate population-level estimates. For both DHS and NSFG data, to achieve adequate sample sizes while ensuring cohorts of women were exposed to broadly similar social conditions, we pool available survey rounds and organize respondents into 10-year birth cohorts. The final cohorts analyzed correspond to women born between 1938–1948 for the United States, and between 1966–1975 for Colombia, the Dominican Republic, and Peru.
@@
Because DHS sample sizes and the number of survey rounds differ widely across countries, we implemented a further filter based on cohort size. We selected DHS countries where the chosen cohort included a minimum of 4,000 women, a threshold established to guarantee reasonably precise estimates of both age-specific total and unplanned fertility rates across the entire reproductive span. This resulted in the inclusion of three DHS countries: Colombia, Peru, and the Dominican Republic. Together with the United States, these four nations constitute our study sample, offering a valuable cross-section of different fertility patterns and socio-economic contexts that allows for a robust examination of our model’s performance across diverse demographic and social settings.
We compute the conventional age-specific fertility rate by dividing the number of births at each age by the number of person-years of exposure contributed at that age. To test how information on birth intentionality improves our estimates, we also construct an age-specific unplanned fertility rate (ASUFRs). This rate is defined as the number of unplanned births at a given age divided by person-years of exposure. The classification of a birth as unplanned is based on a harmonized strategy across the DHS and NSFG that combines direct survey questions about birth timing with a parity check (i.e., whether the birth exceeded the mother’s ideal family size). This hybrid approach is designed to mitigate recall bias and ex-post rationalization, which are known to affect estimates based solely on intention questions (Casterline and El-Zeini, 2007).
@@ 4 Simulation-Based Inference
Inferring parameters for complex process-based models is often challenging due to intractable likelihood functions. Neural Posterior Estimation (NPE) offers a powerful simulation-based inference approach that circumvents these difficulties by directly learning an approximation to the posterior distribution $p(\theta\mid x)$ from simulated data pairs $(\theta,x)$. In practice, NPE proceeds by drawing parameters $\theta$ from a prior distribution ${p}(\theta)$, using the simulator to generate the corresponding data $x\sim p(x\mid\theta)$, and then training a flexible neural density estimator, $q_{\phi}(\theta\mid x)$ (with network parameters $\phi$), on these simulated pairs to approximate the true posterior.
While basic NPE is effective, our study employs a powerful sequential variant known as Automatic Posterior Transformation (APT), sometimes referred to as SNPE-C (Greenberg et al., 2019). Like other sequential neural posterior estimation methods, APT enhances inferential efficiency by drawing parameter samples from an iteratively updated proposal distribution, $\tilde{p}(\theta)$, rather than solely from the fixed prior. The key challenge in this approach is to correct for the use of a proposal distribution, as naively training on the resulting samples would lead to an incorrect posterior estimate. APT solves this by analytically transforming the posterior estimate itself. For each simulated data point drawn from a proposal $\tilde{p}_{r}(\theta)$, it re-weights the target density using the known ratio $\tilde{p}_{r}(\theta)/p(\theta)$. The network is then trained by minimizing the loss with respect to this transformed "proposal posterior" (see line 10 of Algorithm 1). A key advantage of this formulation is that because the loss is valid for any proposal, the network can be trained on the cumulative set of all simulations generated across all previous rounds. This procedure allows the network to learn the true posterior without relying on the potentially unstable importance weights used by other methods.
@@
4.1 Inference Scenarios
Simulation-Based Inference typically relies on summarizing potentially high-dimensional data into lower-dimensional, sufficiently informative statistics. This can be done either through explicit, pre-defined summaries or via learned embeddings. The ability to work with a low-dimensional yet descriptive representation of the data is often essential for making parameter inference tractable, especially for complex simulators.
When working with aggregate data, as in our study, the dimensionality reduction is an inherent feature of the observation process. For instance, vital statistics or administrative registers implicitly aggregate thousands of individual fertility trajectories into relatively low-dimensional summaries like age-specific fertility rates. We thus inherit a pre-summarized dataset, limiting our ability to engineer or learn optimal summary statistics directly from the underlying micro-level process.
A critical question then arises: do these aggregate data provide sufficient information to tightly constrain the posterior distributions of the parameters in our individual-level model? The core challenge is that different combinations of micro-parameters (e.g., those governing desired family size, timing of intentional reproduction, or contraceptive failure) could potentially yield very similar aggregate fertility schedules, resulting in high posterior uncertainty.
While ASFRs represent the baseline data available in most settings, demographic analyses can sometimes draw upon richer sources of information. We therefore test how the parameter estimation process can be strengthened by incorporating such additional information. This can be done in two primary ways: either by augmenting the data with more detailed summary statistics or by encoding external knowledge into the model through more informative priors. Given that our model explicitly simulates the planned and unplanned nature of births, information on birth intentionality is a particularly relevant source of additional data. However, since this type of data is costly to collect and often unavailable, it is valuable to assess whether similar gains in estimation accuracy can be achieved by instead using informative priors on key behavioral parameters.
We therefore evaluate parameter recovery under three increasingly informative inference scenarios:
Scenario 1: ASFRs with Weak Priors: This baseline scenario uses only simulated age-specific fertility rates as summary statistics. All model parameters are assigned weakly informative priors.
Scenario 2: ASFRs with Informative Priors. This scenario tests the impact of incorporating plausible external information. The summary statistics are still limited to ASFRs, but we use narrower, more informative priors for the mean desired family size ($\mu_{d}$), the mean gap to intentional reproduction ($\delta_{r}$), and the mean desired birth spacing ($\mu_{b}$). All other parameters retain their weakly informative priors.
Scenario 3: ASFRs and ASUFRs with Weak Priors. This scenario tests the impact of adding more detailed data. We revert to the weakly informative priors from Scenario 1, but we augment the summary statistics to include both simulated ASFRs and age-specific unplanned fertility rates.
The informative priors used in Scenario 2 simulate a context where a researcher might leverage external information to improve estimates. For the mean desired family size ($\mu_{d}$), we construct the prior directly from the empirical survey distribution for the U.S. case, mimicking a situation where such summary data is readily available. For the remaining two parameters, we simulate the process of knowledge transfer from a data-rich to a data-poor setting by using the posteriors from our most data-intensive setup (Scenario 3) as the informative priors for Scenario 2. This mimics a situation where results from a detailed prior study on a similar cohort/country are used to inform a new analysis where only ASFRs are available.
4.2 Prior Specification
For our baseline scenario (Scenario 1), we selected weakly informative priors for all eleven model parameters to allow the data to drive the inference. The specific parameterization of these priors was chosen to cover a broad yet plausible range of values based on existing demographic literature and the inherent constraints of each parameter.
For strictly positive quantities like means and standard deviations (e.g., $\mu_{s},\sigma_{s}$), we used broad Gamma distributions. For example, the prior for the mean age at sexual initiation ($\mu_{s}$) contains 95% of its probability mass concentrated between 14.4 and 28.9 years. This range comfortably contains existing estimates for the mean age at onset of sexual activity across diverse Western countries and cohorts, which typically falls between 16 and 19 years (Wellings et al., 2006; Kerry L. et al., 2017). Likewise, the prior for mean desired family size ($\mu_{d}$) is centered at 4.5 children, with a 95% probability interval of [1.9, 8.1], again comfortably encompassing the available estimates for the cohorts studied (Westoff and Ryder, 2015; Bongaarts and Lightbourne, 1992; Sobotka and Beaujouan, 2014). For parameters where reliable external estimates are lacking, such as the gap to intentional reproduction ($\delta_{r}$) and mean desired birth spacing ($\mu_{b}$), we employed Uniform priors to define very wide yet plausible ranges (0 to 8 years for $\delta_{r}$ and 10 to 100 months for $\mu_{b}$). The contraceptive failure probability ($\kappa$) was assigned a Beta distribution. Its prior reflects a wide range of plausible contraceptive efficacy levels for our study cohorts, centered around a 20% failure rate but with significant probability assigned to both lower and higher values. Finally, the priors for the Bernstein coefficients ($\beta_{1}$ and $\beta_{2}$) were selected to ensure the resulting age-fecundability curves would generously contain available estimates of conception probabilities at early, peak, and later reproductive ages (Schwartz and Mayaux, 1982; Dunson et al., 2002; Wesselink et al., 2017; ACO, 2014).